# OmniSpatial: Towards Comprehensive Spatial Reasoning Benchmark for Vision Language Models

### v1

@@ Abstract
Spatial reasoning is a key aspect of cognitive psychology and remains a major bottleneck for current vision-language models (VLMs). While extensive research has aimed to evaluate or improve VLMs’ understanding of basic spatial relations, such as distinguishing left from right, near from far, and object counting, these tasks represent only the most fundamental level of spatial reasoning. In this work, we introduce OmniSpatial, a comprehensive and challenging benchmark for spatial reasoning, grounded in cognitive psychology. OmniSpatial covers four major categories: dynamic reasoning, complex spatial logic, spatial interaction, and perspective-taking, with 50 fine-grained subcategories. Through Internet data crawling and careful manual annotation, we construct over 1.5K question-answer pairs. Extensive experiments show that both open- and closed-source VLMs, as well as existing reasoning and spatial understanding models, exhibit significant limitations in comprehensive spatial understanding. We further analyze failure cases and propose potential directions for future research.
@@ 1 Introduction
Spatial reasoning plays a crucial role in bridging visual observation to robotic action [43, 88, 89], autonomous driving, and AR/VR. For models to execute tasks effectively, they must understand spatial relationships to determine appropriate actions. To enhance the spatial understanding of Vision-Language Models (VLMs), prior works [15, 17, 11, 68, 102, 123, 121, 89] have integrated spatial information into datasets, enabling basic forms of spatial reasoning. Various benchmarks [26, 104, 101, 48, 34, 102, 121, 89] have been introduced to systematically evaluate such capabilities, focusing on tasks like recognizing left and right, estimating depth, and constructing cognitive maps [121, 106, 72]. Additionally, spatial reasoning has been applied to manipulation tasks [25, 8, 131, 20, 123], allowing systems to position objects according to specified spatial rules [89].
@@
Despite these advances, existing approaches largely focus on basic spatial understanding, which typically includes identifying simple spatial relations such as relative positions (left, right, front, back), proximity (near, far), and object-level counting. Recent models, including state-of-the-art reasoning systems like ChatGPT o3 [79] and Gemini-2.5-Pro [97], have achieved over 90% accuracy on benchmarks such as SpatialBot-Bench [11] and EmbSpatial [26], suggesting that these basic tasks are approaching saturation, as shown in OmniSpatial: Towards Comprehensive Spatial Reasoning Benchmark for Vision Language Models.
@@
However, complex spatial reasoning remains a significant challenge [35, 5, 83, 50, 82, 16]. Human interaction with the physical world often involves interpreting ambiguous, dynamic, and context-dependent spatial relationships [7, 109, 95]. For example, in an emergency, knowing that an AED is “to the right of the door” is insufficient without understanding schematic diagrams, correlating maps with real-world environments, and planning an efficient route. Similarly, tasks like inserting a knife into a rack or flattening a box demand reasoning about object rotation, deformation, and spatial compatibility—far beyond static object placement.
From the perspective of cognitive psychology, complex spatial reasoning extends beyond basic relations and includes dynamic world knowledge reasoning, interactive spatial behavior with environments or agents, logical analysis of 3D spatial structures, and perspective-taking abilities.
Therefore, to address these limitations and systematically assess higher-order spatial reasoning, we propose OmniSpatial, a comprehensive benchmark capturing the breadth and depth of spatial cognition. OmniSpatial categorizes spatial reasoning into four key dimensions: dynamic reasoning, complex spatial logic, spatial interaction, and perspective-taking, providing a foundation for advancing the next generation of spatially-aware & physics-aware AI systems.
@@
EmbSpatial-Bench [26] 6 Indoor (ScanNet, etc.) Template 2.2K 3.6K
Space3D-Bench [104] 6 Indoor (Replica) Manual 211 1K
Visual Spatial [61] 7 MSCOCO Template 10K 10K
SpatialRGPT-Bench [17] 12 Urban, Indoor, Sim Template 1.4K 1.4K
BLINK-Spatial [34] 14 MSCOCO Manual 286 286
What’s up [48] 6 Household, GQA, COCO Template 5K 5K
Spatial-MM [101] 4 Internet Template 2.3K 2.3K
RoboSpatial [102] 4 Indoor, tabletop Template 1M 3M
SpatialVLM [15] 2 WebLi Template 546 546
SpatialBot-Bench [11] 5 COCO,VG,RTX Manual 200 360
VSI-Bench [121] 8 Indoor Template 288 5K
OmniSpatial (Ours) 50 Internet Manual 1387 1.5K
Table 1: Comparison with other spatial reasoning benchmarks. A comparison between OmniSpatial and other existing spatial reasoning benchmarks reveals that, under the inclusion of Embodied Frames Compatibility, OmniSpatial avoids template-based annotations, features highly diverse data, and includes a significantly larger number of tasks.
@@ 2 Preliminaries: Visual–Spatial Reasoning
Spatial reasoning constitutes the cognitive bridge between visual perception and action planning in embodied agents. We define visual–spatial reasoning as the capacity of an artificial system to infer, predict, and manipulate spatial properties of the world from visual observations. Formally, let an RGB observation stream be $\mathbf{I}_{1:T}$ and a task-specific query be $q$. A model possesses visual–spatial reasoning if it can learn a mapping
$f:(\mathbf{I}_{1:T},\,q)\;\longrightarrow\;a,$
where $a$ belongs to a well-defined action or answer space whose correctness can be verified in the physical or simulated environment. This definition excludes non-visual priors so that improvements can be attributed to visual reasoning itself, yet it remains compatible with multi-modal extensions discussed in §1.
@@ 2.1 Taxonomy of Visual–Spatial Reasoning
Our classification stems from two complementary motivations. (i) Cognitive-psychology foundations. Decades of research on spatial cognition isolate partly independent faculties—e.g., spatial visualization, mental rotation, perspective taking, and spatial updating—that can be psychometrically evaluated in humans [13, 71]. These constructs offer a principled scaffold for analysing how agents perceive, reason about, and act within space. (ii) The need to move beyond basic spatial relations. Current spatial benchmarks are nearly saturated on tasks such as left–right discrimination, front–back identification, and object counting [15, 11, 17, 26, 104]. Real-world embodied tasks, however, require richer reasoning [50, 82, 16, 103, 53] about scene dynamics, multi-step spatial logic, physically grounded interaction, and viewpoint transformation.
Guided by these two strands, we partition spatial reasoning into four orthogonal yet complementary dimensions—dynamic reasoning, complex spatial logic, spatial interaction, and perspective taking. Each dimension aligns with a specific cognitive faculty and targets a distinct class of spatial challenges that is under-represented in prior work, enabling us to construct benchmarks that probe the full spectrum of complex spatial cognition while remaining anchored in established psychological theory.
Dynamic Reasoning involves predicting motion and temporal changes in spatial environments. This capability is crucial for tasks such as robotic control and autonomous navigation, where anticipating object trajectories ensures adaptive decision-making in dynamic contexts.
Complex Spatial Logic encompasses high-level reasoning about spatial relationships, transformations, and geometric structures. It is fundamental to problem-solving in domains such as engineering, design, and robotic manipulation, where understanding spatial configurations and predicting structural changes are essential.
Spatial Interaction focuses on engagement with spatial environments, where reasoning is guided by environmental constraints and task objectives. This includes applications such as path planning, obstacle avoidance, and real-time decision-making based on spatial feedback.
Perspective Taking refers to the ability to adopt different viewpoints, which is crucial for navigation, social cognition, and spatial awareness. It allows for understanding spatial relationships from multiple perspectives, facilitating decision-making in multi-agent environments and adaptive problem-solving.
@@ 2.2 Rationale for Classification
This taxonomy is designed to be both theoretically comprehensive and practically applicable. Dynamic reasoning emphasizes motion prediction, which is critical for navigation and interaction. Complex spatial logic provides a foundation for advanced cognitive tasks involving transformation and abstraction. Spatial interaction highlights real-time engagement with environments, while perspective taking captures cognitive flexibility in spatial reasoning. By structuring visual-spatial reasoning in this manner, we establish a framework that enhances the evaluation of spatial cognition models and their applications in artificial intelligence, robotics, and human cognition research.
@@
3 OmniSpatial: Towards Comprehensive Spatial Reasoning Benchmark
3.1 Overview
We introduce OmniSpatial, a benchmark for the quantitative evaluation of VLMs in visual-spatial reasoning. Unlike large-scale benchmarks focused on a few tasks, OmniSpatial prioritizes task diversity and structured categorization over sheer dataset size. It consists of 1533 high-quality question-answer pairs, designed to challenge models beyond simple pattern matching or statistical learning, ensuring that spatial reasoning capabilities are tested rigorously, even with limited data.
Our dataset is curated from web searches, standardized tests, driving exam questions, and existing datasets including MME [57], HOI4D [67], etc.This diverse sourcing enhances the realism, complexity, and cross-domain generalization of the tasks. Web-sourced images cover natural environments, architecture, and daily life, adding visual complexity. Standardized tests from psychology and cognitive science provide scientifically rigorous spatial reasoning challenges. Driving exam questions introduce real-world dynamic interactions, such as road rule comprehension and motion prediction. Additionally, datasets like HOI4D contribute multi-resolution, varying illumination, and viewpoint diversity, as well as embodied intelligence tasks involving human-object interactions.
To ensure comprehensive evaluation, we categorize tasks into 4 major spatial reasoning types, further divided into 50 fine-grained task categories, as shown in Figs. 6 and 2. For example, perspective-taking tasks range from simple spatial judgments (“Is the opponent’s weapon to the left or right of the blue athlete?”) to complex motion prediction (“The blue athlete is blocking an incoming attack from the red athlete—what will happen next?”). Spatial interaction tasks include both static collision assessments (“Is the vehicle too close to the one in front?”) and dynamic environment reasoning (“A car door ahead is opening—a pedestrian exiting?”). This multi-level task structure ensures models are tested on diverse and progressively complex spatial reasoning challenges.
We emphasize precise image-task alignment in both dataset selection and annotation. Each question-answer pair is manually curated and reviewed in multiple rounds to ensure accuracy, consistency, and minimal ambiguity. Unlike automated annotation, manual refinement guarantees a high correlation between visual scenes and spatial reasoning tasks, strengthening OmniSpatial’s reliability as a benchmark for future research.
@@ 3.2 Benchmark Construction
This section describes how we collect data and annotations for OmniSpatial, as illustrated in Fig. 3.
@@ 3.2.1 Data Collection
As outlined in Section 2, our study defines 4 spatial categories with corresponding task types. To ensure a diverse and information-rich dataset, we design targeted search strategies for each task type, optimizing data collection for relevance, diversity, and complexity.
Web Images Web images constitute a major portion of our dataset, particularly for Perspective Taking, Dynamic Reasoning, and Spatial Interaction tasks. To ensure relevance and diversity, we formulate task-specific search terms reflecting distinct spatial scenarios, such as “indoor layout” and “furniture arrangement” for spatial relationship tasks. To minimize AI-generated image interference, we append filters like “-ai” and “-generated” to search queries. Using Google’s Custom Search JSON API, Web Robotic Process Automation (Web RPA), and manual searches, we retrieve thousands of images efficiently. A rigorous screening process is followed to eliminate irrelevant, low-resolution, or overly compressed images. We also removed images lacking spatial complexity, such as isolated static objects in Spatial Interaction tasks. The final dataset balances visual complexity and realism, enhancing its generalization capability across multiple tasks and scenarios while ensuring high-quality training data for VLMs. All the images are under MIT or CC-BY 4.0 licence.
Exam-Based Test Questions  Designed to assess complex spatial reasoning, these questions emphasize abstract spatial logic, including 3D transformations, rotations, and perspective shifts. We collect a wide range of publicly available spatial cognition tests via web scraping and manual curation. Noting variations in task focus, difficulty, and reasoning style, we categorized the questions to maintain a balanced representation. To prevent overrepresentation of any single reasoning type, we curate the dataset to ensure diverse question formats. Redundant, overly similar, or knowledge-intensive questions are removed, prioritizing those that purely test spatial reasoning. This refinement process significantly enhances the dataset’s quality and challenge level, providing a robust evaluation framework for VLMs.
@@
Driving Test Questions  These questions evaluate spatial interaction in dynamic environments, focusing on distance estimation, right-of-way rules, and hazard detection. We source them from three primary channels: (i) Image-based multiple-choice questions from driving test websites in at least three countries, capturing diverse traffic rules and scenarios. (ii) Online question banks covering standardized spatial judgment tasks such as turning, lane changes, and parking decisions. (iii) Interactive driving videos from U.S. test resources, where we extract frames, annotate bounding boxes, and formulate contextual questions (“Which bounding box highlights a potential traffic risk?”). This multi-source approach ensures a realistic and challenging dataset, enhancing VLM adaptability in traffic-related spatial reasoning.
@@
Existing Datasets  We integrate two key datasets: MME [57] and HOI4D [67]. MME provides RGB-D data, allowing depth-based spatial inference. We leverage its depth information for physics-based questions such as “If a red car passes me in 5 seconds, what speed should I maintain?” This ensures realistic distance and motion-based reasoning. HOI4D contains extensive human-object interaction videos. We extract sequential frames to create motion prediction tasks, such as “Where will the hand holding the kettle move next?” By incorporating these datasets, we introduce real-world motion and interaction complexities, further strengthening VLMs’ dynamic reasoning capabilities.

### v2

@@ Abstract
Spatial reasoning is a key aspect of cognitive psychology and remains a bottleneck for current vision-language models (VLMs). While extensive research has aimed to evaluate or improve VLMs’ understanding of basic spatial relations, such as distinguishing left from right, near from far, and object counting, these tasks cover only the most elementary layer of spatial reasoning and are largely approaching saturation in the latest reasoning models. In this work, we introduce OmniSpatial, a comprehensive and challenging benchmark for spatial reasoning, grounded in cognitive psychology. OmniSpatial covers four major categories: dynamic reasoning, complex spatial logic, spatial interaction, and perspective-taking, with 50 fine-grained subcategories. Through careful manual annotation, we construct over 8.4K question-answer pairs. Extensive experiments show that both open- and closed-source VLMs exhibit significant limitations in comprehensive spatial reasoning. We also explore two strategies—PointGraph (explicit scene graph cues) and SpatialCoT (novel-view chain-of-thought)—to bolster spatial reasoning.
@@ 1 Introduction
Spatial reasoning plays a crucial role in bridging visual observation to robotic action (Huang et al., 2024b, Qi et al., 2024; 2025), autonomous driving, and AR/VR. For models to execute tasks effectively, they must understand spatial relationships to determine appropriate actions. To enhance the spatial understanding of Vision-Language Models, prior works (Chen et al., 2024, Cheng et al., 2024, Cai et al., 2024, Ma et al., 2024, Song et al., 2024, Yuan et al., 2024, Yang et al., 2024, Qi et al., 2025) have integrated spatial information into datasets, enabling basic forms of spatial reasoning. Various benchmarks (Du et al., 2024, Szymanska et al., 2024, Shiri et al., 2024, Fu et al., 2024b, Song et al., 2024, Yang et al., 2024) have been introduced to systematically evaluate such capabilities, focusing on tasks like recognizing left and right, estimating depth, and constructing cognitive maps (Yang et al., 2024, Tolman, 1948, Momennejad et al., 2023). Additionally, spatial reasoning has been applied to manipulation tasks (Driess et al., 2023, Qi et al., 2025, Yuan et al., 2024), allowing systems to position objects according to specified spatial rules.
@@
However, existing benchmarks still target basic spatial understanding, such as position relationship (left, right, front, back), proximity (near, far), and object counting. We use the latest reasoning models and agents to evaluate these benchmarks, such as o3 (OpenAI, 2025b) and Gemini-2.5-Pro (Reid et al., 2024). The results are shown in the lower left of OmniSpatial: Towards Comprehensive Spatial Reasoning Benchmark for Vision Language Models. These models have achieved over 90% accuracy on previous benchmarks such as SpatialBot-Bench (Cai et al., 2024) and EmbSpatial (Du et al., 2024), suggesting that these basic tasks are approaching saturation.
@@
We believe complex spatial reasoning remains a significant challenge (Gardner, 2011, Baddeley, 1998, Previc, 1998, Kosoy et al., 2025, Pothiraj et al., 2025, Chen et al., 2025). Human interaction with the physical world often involves interpreting ambiguous, dynamic, and context-dependent spatial relationships (Bar-Anan et al., 2006, Trope & Liberman, 2010, Ramalho et al., 2018). For example, in an emergency, knowing that an AED is “to the right of the door” is insufficient without understanding schematic diagrams, correlating maps with real-world environments, and planning an efficient route. Similarly, tasks like inserting a knife into a rack or flattening a box demand reasoning about object rotation, deformation, and spatial compatibility—far beyond static object placement.
From the perspective of cognitive psychology, complex spatial reasoning goes beyond simple relational judgments, encompassing dynamic world-knowledge reasoning, interactive spatial behavior with environments or agents, logical analysis of three-dimensional structures, and perspective-taking abilities. Motivated by these challenges, we introduce OmniSpatial, a comprehensive benchmark designed to capture the breadth and depth of spatial cognition. OmniSpatial systematically categorizes spatial reasoning into four core dimensions—dynamic reasoning, complex spatial logic, spatial interaction, and perspective-taking—thus providing a principled foundation for developing next-generation spatially- and physically-aware AI systems.
Table 1: Comparison with other spatial reasoning benchmarks. A comparison between OmniSpatial and other existing spatial reasoning benchmarks. OmniSpatial avoids template-based annotations, features highly diverse data, and includes a significantly larger number of tasks.
@@
EmbSpatial-Bench (Du et al., 2024) 6 Indoor (ScanNet, etc.) Template 2.2K 3.6K
Space3D-Bench (Szymanska et al., 2024) 6 Indoor (Replica) Manual 211 1K
Visual Spatial (Liu et al., 2023a) 7 MSCOCO Template 10K 10K
SpatialRGPT-Bench (Cheng et al., 2024) 12 Urban, Indoor, Sim Template 1.4K 1.4K
What’s up (Kamath et al., 2023) 6 Household, GQA, COCO Template 5K 5K
Spatial-MM (Shiri et al., 2024) 4 Internet Template 2.3K 2.3K
RoboSpatial (Song et al., 2024) 4 Indoor, tabletop Template 1M 3M
SpatialVLM (Chen et al., 2024) 2 WebLi Template 546 546
SpatialBot-Bench (Cai et al., 2024) 5 COCO,VG,RTX Manual 200 360
VSI-Bench (Yang et al., 2024) 8 Indoor Template 288 5K
OmniSpatial (Ours) 50 Internet Manual 6.5K 8.4K
@@ 2 Preliminaries: Visual–Spatial Reasoning
Spatial reasoning constitutes the cognitive bridge between visual perception and geometric understanding. We define visual–spatial reasoning as the capacity of an artificial system to infer, predict, and reason spatial properties of the world from visual observations. Formally, let an RGB observation stream be $\mathbf{I}_{1:T}$ and a task‑specific query be $q$. A model possesses visual–spatial reasoning that learn a mapping:
$f:(\mathbf{I}_{1:T},\,q)\;\longrightarrow\;a,$ (1)
where $a$ belongs to a well‑defined action or answer space whose correctness can be verified in the physical or simulated environment. This definition excludes non‑visual priors so that improvements can be attributed to visual reasoning itself, yet it remains compatible with multi‑modal extensions discussed in §1.
@@ 2.1 Taxonomy of Visual–Spatial Reasoning
Our taxonomy is motivated by two complementary perspectives. (i) Cognitive psychology foundations. Prior research on spatial cognition highlights partly independent faculties—such as visualization, mental rotation, perspective taking, and spatial updating—that can be systematically assessed in humans (Chabris et al., 2006, Meneghetti et al., 2022). These constructs provide a principled scaffold for analyzing how agents perceive, reason about, and act within space. (ii) Moving beyond basic spatial relations. Existing benchmarks are nearly saturated on simple tasks like left–right discrimination, front–back identification, and object counting (Chen et al., 2024, Cai et al., 2024, Du et al., 2024, Szymanska et al., 2024). Yet real-world embodied tasks demand richer reasoning (Kosoy et al., 2025, Pothiraj et al., 2025, Chen et al., 2025, Stogiannidis et al., 2025, Lee et al., 2025) about scene dynamics, multi-step logic, physical interaction, and viewpoint transformation.
Guided by these considerations, we partition visual–spatial reasoning into four complementary dimensions: dynamic reasoning, complex spatial logic, spatial interaction, and perspective taking. Each dimension corresponds to a specific cognitive faculty and targets under-explored challenges in prior work, enabling us to probe a broader spectrum of spatial cognition while remaining grounded in psychological theory.
Dynamic Reasoning concerns inferring motion and temporal change from visual evidence. While our benchmark primarily uses static or sparsely sampled frames, such inference is crucial for adaptive decision-making in domains like robotics and navigation.
Complex Spatial Logic involves higher-order reasoning about relations, transformations, and geometric structures. It underpins problem-solving in design, engineering, and manipulation, where anticipating structural or relational changes is essential.
Spatial Interaction emphasizes reasoning guided by environmental constraints and task goals, covering skills such as path planning, obstacle avoidance, and context-aware action selection.
Perspective Taking captures the ability to adopt alternative viewpoints, supporting navigation, social cognition, and multi-agent coordination. It enables understanding relations from diverse perspectives and fosters flexible problem-solving.
@@ 2.2 Rationale for Classification
This taxonomy balances theoretical comprehensiveness with practical applicability. Dynamic reasoning highlights motion inference, complex logic captures abstract transformations, spatial interaction addresses real-time engagement, and perspective taking reflects cognitive flexibility. Together, these dimensions provide a framework for evaluating spatial reasoning in AI, robotics, and human cognition.
@@
3 OmniSpatial: Comprehensive Spatial Reasoning Benchmark
3.1 Overview
We present OmniSpatial, a comprehensive benchmark designed to evaluate vision-language models on spatial reasoning. Rather than pursuing sheer data volume, OmniSpatial emphasizes diversity, structure, and rigor. It now consists of 8.4K carefully curated question–answer pairs, substantially larger than earlier prototypes, and covers a broad spectrum of scenarios that demand reasoning beyond pattern recognition.
The dataset integrates heterogeneous sources—web imagery, standardized cognitive tests, driving-exam questions, and prior dataset images such as MME (Liang et al., 2024) and HOI4D (Liu et al., 2022). This mixture enriches both realism and complexity: natural images capture everyday environments and architectures; psychology-inspired tasks introduce scientifically grounded challenges; driving exams provide safety-critical dynamic reasoning; and embodied datasets contribute varied resolutions, viewpoints, and human–object interactions.
Tasks are organized into 4 major categories and 50 fine-grained subtypes as shown in Figs.˜6 and 2, spanning from basic perspective-taking to dynamic motion prediction and spatial interaction in cluttered scenes. Each item is manually designed and reviewed through multi-round annotation, ensuring accuracy, consistency, and minimal ambiguity. This careful curation yields a benchmark that not only broadens the coverage of spatial reasoning but also establishes a reliable ground for evaluating and advancing future multimodal intelligence.
@@ 3.2.1 Data Collection
As described in Section˜2, we define four spatial categories and corresponding task types. To build a diverse and information-rich dataset, we design targeted search strategies for each type, optimizing for relevance, diversity, and complexity.
Web Images. Web images form a major part of our data, especially for perspective-taking, dynamic reasoning, and spatial interaction. We use task-specific search terms (e.g., “indoor layout,” “furniture arrangement”) and append filters (“-ai,” “-generated”) to reduce synthetic content. Images are retrieved via Google’s Custom Search JSON API, Web RPA, and manual search, followed by strict filtering to remove irrelevant, low-resolution, or spatially trivial cases (e.g., isolated static objects). The resulting set balances realism and complexity, ensuring broad task coverage. All images are under MIT or CC-BY 4.0 license.
Exam-Based Test Questions. To capture abstract spatial logic, such as 3D transformations, rotations, and perspective shifts, we collect public spatial cognition tests through web scraping and manual curation. We categorize questions by focus and difficulty to maintain balance, removing redundant or knowledge-heavy items in favor of those targeting pure spatial reasoning. This refinement increases challenge diversity and improves benchmark quality.
@@
Driving Test Questions. To evaluate reasoning in dynamic environments, we source tasks from three channels: (i) image-based multiple-choice questions from driving exam websites across at least three countries, (ii) online banks of standardized tasks like turning, lane changing, and parking, and (iii) interactive U.S. driving test videos, from which we extract frames, annotate bounding boxes, and design contextual queries (e.g., “Which bounding box indicates a potential traffic hazard?”). This combination yields realistic and challenging traffic scenarios, enhancing VLM adaptability to safety-critical reasoning tasks.
@@
Existing Dataset Images  We integrate two key data sources: MME (Liang et al., 2024) and HOI4D (Liu et al., 2022). MME provides RGB-D data, allowing depth-based spatial inference. We leverage its depth information and manually propose physics-based questions such as “If a red car passes me in 5 seconds, what speed should I maintain?” This ensures realistic distance and motion-based reasoning. HOI4D contains extensive human-object interaction videos. We extract sequential frames to create motion prediction tasks, such as “Where will the hand holding the kettle move next?” By incorporating these datasets, we introduce real-world motion and interaction complexities, further strengthening VLMs’ dynamic reasoning capabilities. Furthermore, to extend our dataset to model training, we partition the dataset into a 1.5K test set and a 6.9K training set. The test set is entirely human-annotated and further integrates samples from several existing datasets. including SpatialViz  (Wang et al., 2025), PhysBench (Chow et al., 2025), ViewSpatial (Li et al., 2025), and DrivingVQA (Corbière et al., 2025), which substantially enhance the diversity of our dataset.